Citation Builders, Inc., builds office buildings and single-family homes. The office buildings are constructed under contract with reputable buyers. The homes are constructed in developments ranging from 10−20 homes and are typically sold during construction or soon after. To secure the home upon completion, buyers must pay a deposit of 10% of the price of the home with the remaining balance due upon completion of the house and transfer of title. Failure to pay the full amount results in forfeiture of the down payment. Occasionally, homes remain unsold for as long as three months after construction. In these situations, sales price reductions are used to promote the sale. During 2021, Citation began construction of an office building for Altamont Corporation. The total contract price is $25 million. Costs incurred, estimated costs to complete at year-end, billings, and cash collections for the life of the contract are as follows: 2021 2022 2023 Costs incurred during the year $ 5,000,000 $ 11,875,000 $ 5,625,000 Estimated costs to complete as of year-end 15,000,000 5,625,000 − Billings during the year 2,500,000 12,500,000 10,000,000 Cash collections during the year 2,250,000 11,150,000 11,600,000 Also during 2021, Citation began a development consisting of 12 identical homes. Citation estimated that each home will sell for $900,000, but individual sales prices are negotiated with buyers. Deposits were received for eight of the homes, three of which were completed during 2021 and paid for in full for $900,000 each by the buyers. The completed homes cost $675,000 each to construct. The construction costs incurred during 2021 for the nine uncompleted homes totaled $4,050,000.

Required: 1. Which method is most equivalent to recognizing revenue at the point of delivery? 2. Answer the following questions assuming that Citation concludes it does not qualify for revenue recognition over time for its office building contracts: 2-a. How much revenue related to this contract will Citation report in its 2021 and 2022 income statements? 2-b. What is the amount of gross profit or loss to be recognized for the Altamont contract during 2021 and 2022? 2-c. What will Citation report in its December 31, 2021, balance sheet related to this contract? (Ignore cash.) 3. Answer the following questions assuming that Citation recognizes revenue over time according to percentage of completion for its office building contracts. 3-a. How much revenue related to this contract will Citation report in its 2021 and 2022 income statements? 3-b. What is the amount of gross profit or loss to be recognized for the Altamont contract during 2021 and 2022? 3-c. What will Citation report in its December 31, 2021, balance sheet related to this contract? (Ignore cash.) 4. Assume the same information for 2021 and 2022, but that as of year-end 2022 the estimated cost to complete the office building is $11,250,000. Citation recognizes revenue over time according to percentage of completion for its office building contracts. 4-a. How much revenue related to this contract will Citation report in the 2022 income statement? 4-b. What is the amount of gross profit or loss to be recognized for the Altamont contract during 2022? 4-c. What will Citation report in its 2022 balance sheet related to this contract? (Ignore cash.) 5. Which method of accounting should Citation Builders, Inc adopt for its single-family houses? 6. What will Citation report in its 2021 income statement and 2021 balance sheet related to the single-family home business (ignore cash in the balance sheet)?
Answer #1

Dear Student,

As per the HomeworkLib policy, only first four parts. Kindly take note ofit

Part 1

Completed contract

When a long term construction contract is completed, the revenue recognized at completion should be equal to revenue recognized at the delivery point of time. To recognize revenue over the period of time requires to allocating expected revenues and costs to each construction period. The allocation basis is determine on the basis of each period’s project cost incurred in proportion to the total estimated cost of project.

Part 2

2021

2022

Contract price

25000000

25000000

Actual costs to date

5,000,000

16875000

(5000000+11875000)

Estimated costs to complete

15000000

5625000

Total estimated costs

20000000

22500000

Estimated gross profit

5000000

2500000

A. Revenue recognition:

2021 : $0

2022 : $0

total Revenues will be recognized only in 2020

B. Gross profit recognition:

2021 : $0

2022 : $0

gross profit will be recognized only in 2023

C. Citation will report in its December 31, 2021, balance sheet related to this contract = $2500000 (Costs – billings = 5000000-2500000 = 2500000)

Balance sheet (partial)

At December 31, 2021

Current assets:

Accounts receivable (2500000-2250000)

250000

Costs in excess of billings

2500000

Current liabilities:

Part 3

2021

2022

Contract price

25000000

25000000

Actual costs to date

5,000,000

16875000

(5000000+11875000)

Estimated costs to complete

15000000

5625000

Total estimated costs

20000000

22500000

Estimated gross profit

5000000

2500000

A Revenue recognition:

2021

Revenue: 5000000/20000000 = 25%*25000000 = 6250000

2022

Revenue: 16875000/22500000 = 75%*25000000 = 18750000-6250000 = 12500000

B. Gross profit recognition:

2021 = 6250000-5000000 = 1250000

2022 = 12500000-11875000 = 625000

C . Citation will report in its December 31, 2021, balance sheet related to this contract = $3750000 (Costs + profit – billings = 5000000+1250000-2500000 = 3750000)

Current assets:

Accounts receivable

250000

Costs and profit in excess of billings

3750000

Current liabilities:

Accounts receivable = billings – cash collection = 2500000-2250000 = 250000

Part 4 A & B

Revenue reported

$7000000

Gross profit or loss to be recognized

$(3500000)

2021

2022

Contract price

25000000

25000000

Actual costs to date

5,000,000

16875000

(5000000+11875000)

Estimated costs to complete

15000000

11250000

Total estimated costs

20000000

28125000

Estimated gross profit

5000000

(3125000)

A. Revenue recognition:

Total revenue recognized to date= (percentage complete) ×(total revenue)

= (16875000/28125000) x 25000000 = 15000000

Revenue recognized in 2022= total – revenue recognized in prior periods = 15000000-6250000 = $8750000

(Revenue recognized in prior period (2021) = 5000000/20000000*25000000 = 6250000)

B. Gross profit or loss recognition:

Gross loss = overall loss – previously recognized gross profit = (3125000)-1250000 = (4375000)

(previously recognized gross profit (2021) = 6250000-5000000=1250000)

Part 4 C

Current assets:

Accounts receivable

2600000

Current liabilities:

Billings in excess of Costs and profit

1250000

Accounts receivable = (2500000-1250000)+(12500000-11150000) = 2600000

Citation will report in its December 31, 2019, balance sheet related to this contract = billings – (cost + profit ) = (2500000+12500000)-(5000000+11875000+1250000-4375000) = 1250000